| ID | 
Date | 
Author | 
Category | 
Subject | 
| 
  130  
 | 
Thu Apr  3 18:53:02 2025 | 
Usama, Kathrin | General | Screenshots | 
 | 
| 
  129  
 | 
Thu Apr  3 18:01:49 2025 | 
Ruijiu, Carlo | Analysis | Identification of 188Hf | 
 | 
| 
  128  
 | 
Thu Apr  3 17:49:57 2025 | 
Usama, Kathrin | General | Screenshots | 
 | 
| 
  127  
 | 
Thu Apr  3 17:38:47 2025 | 
Shahab, Zac, Carlo, Esther and many more | DAQ | RSAs run again! | 
first file name RSA01....17.38 onwards.  | 
| 
  126  
 | 
Thu Apr  3 17:01:07 2025 | 
CJ Griffin, M Cockshutt, G Leckenby | General | Moving PLEIADES detectors in | 
The inner PLEIADES is moved into position close to the beam.
Device control for the outside pocket is not currently working so that pocket remains in its outward position.
NTCAP/Schottky data acquisition stopped during pocket position tests.
Inner detector (x86l-124 BGO)
Device: GE01DD2IG
Position: +20 mm
Scraping beam: not possible
Outer detector (x86l-131 GAGG)
Device: GE01DD2AG
Position: -127 mm
Scraping beam:   | 
| 
  125  
 | 
Thu Apr  3 16:44:18 2025 | 
Usama, Kathrin | General | Screenshots | 
 | 
| 
  124  
 | 
Thu Apr  3 15:57:16 2025 | 
Shahab, Yue, Zac, Esther | DAQ | Data acquisitions running | 
All data acquisitions (both RSAs and NTCAP) are now running.  | 
| 
  123  
 | 
Thu Apr  3 15:55:04 2025 | 
Shahab, Yue, Esther | Calibration | Schottky settings | 
Settings similar to the 2-photon experiment.
Central frequency: 411.742243 MHz (based on Ruijius identification for 188Hf72+)  | 
| 
  122  
 | 
Thu Apr  3 15:21:42 2025 | 
CJG | General | PLEIADES shift instructions | 
PLEIADES shift instructions available at:
https://docs.google.com/document/d/19b4wyugC3OpVAjwt1ONM9u8EYdMJsBpL7GVNzoziv0M/edit?usp=sharing
Leakage current tracking:
https://docs.google.com/spreadsheets/d/1Gt1tFHwn4XFrJ5_-nQvTi635UD0aHgN6gCdXDTHQdkI/edit?gid=299353113#gid=299353113
Leakage currents should be documented every ~1-2 hours while the detectors are biased.  | 
| 
  121  
 | 
Thu Apr  3 13:36:17 2025 | 
Ruijiu,Yuri, Wolfram,Zsolt | Analysis | 188Hf identification | 
 | 
| 
  120  
 | 
Thu Apr  3 13:32:42 2025 | 
CN, ID | General | Screenshots | 
 | 
| 
  119  
 | 
Thu Apr  3 11:51:00 2025 | 
CN, ID | DAQ | RSA started for 188Hf ESR tuning | 
Started both RSAs for the tuning of the ESR to 188Hf. Both pockets are in parking position OUTSIDE.  | 
| 
  118  
 | 
Thu Apr  3 10:07:34 2025 | 
CN, ZP, CJG, MC | General | Screenshots | 
 | 
| 
  117  
 | 
Thu Apr  3 09:53:02 2025 | 
Ruijiu Chen,Wolfram | Analysis | Good injection ratio | 
Currently, we only use 37% of the injections in our data analysis. However, 63% of the injections have a height below 100, and these may include 
single ion injections. We need to visually inspect these injections to identify the single ion injections.  | 
| 
  116  
 | 
Thu Apr  3 09:43:14 2025 | 
All | Accelerator | Stop 194Pb, switch to 208Pb82+ | 
The repair of SIS electron cooler power supplies
In the meantime we received a few injections and confirmed that 194Pb setting is still as it was before entering into the ESR
Now we change to 208Pb82+ to do several ESR accelerator tests
Esis = 517 MeV/u
Change the beamline to 208 Pb 82+ and the ESR to 208 b 82+ and NO other changes  | 
| 
  115  
 | 
Thu Apr  3 09:34:42 2025 | 
ID | General | Finished in the ESR as of 9:05 | 
We had stopped the RSAs at 8:00 (call for restart)
Time stamp computer stopped and started at 9:35  | 
| 
  114  
 | 
Thu Apr  3 08:02:04 2025 | 
Team PLEIADES (JG, ID, CJG, MC) | Detectors | Entering ESR for PLEIADES | 
We are going to enter the ESR to check on the PLEIADES pocket detectors and install a nitrogen cooling.  | 
| 
  113  
 | 
Thu Apr  3 07:59:09 2025 | 
ZP, CN, ID | General | Screenshots | 
Screenshots before opening ESR.  | 
| 
  112  
 | 
Thu Apr  3 06:56:31 2025 | 
JEL, ZN | Runs | Screenshots | 
 | 
| 
  111  
 | 
Thu Apr  3 06:33:08 2025 | 
Zac, Johan | General | Beam lost at 06:27, returned at 06:34 | 
Spike in beam intensity at 06:27 (5x usual intensity) and then beam is lost. HKR says it's a problem with the magnet's power supply. Once the beam returned, the peaks are shifted to the left but are slowly shifting back into position.  |